OpenStudy (anonymous):

Simplify 5x^3/7x^3+x^4

OpenStudy (anonymous):

\[\frac{ 5x^3 }{ 7x^3+x^4 }\]

OpenStudy (anonymous):

okay so what do you think we do first? @Alice2112

OpenStudy (anonymous):

I have no idea I have 2 days and like 20 assiments

OpenStudy (tkhunny):

You CAN'T have NO idea. Find all the common factors. That's all.

OpenStudy (anonymous):

5 and 7?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

\[\frac{ 5 }{ 7+x } where X \neq 0,-7\]

OpenStudy (anonymous):

\[\frac{ 5 }{ 7+x } where X \neq -7\]

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

What is the question? @Alice2112

OpenStudy (anonymous):

Can you help im not sure what the answer is i need to simplify \[\frac{ 5x^3 }{ 7x^3+x^4 }\]

Directrix (directrix):

ok, just a second.

Directrix (directrix):

Do you agree with this so far?